mormaki83:

Identify the slope and y-intercept of the line whose equation is given. Write the y-intercept as an ordered pair. (Remember the slope-intercept form is 4072-03-01-03-00_files/i0150000.jpg) q = 2 - r a. slope = -2; y-intercept (0, 0) b. slope = 2; y-intercept (0, -1) c. slope = -1; y-intercept (0, 2) d. *question has been edited by a moderator to supply missing information*

Mercury:

treating r as the independent variable and q as the dependent variable we can re-write the equation as q = -r + 2 which, compared to y = mx + b, what are your m and b values?